The length of mercury column containing mercury in a glass thermometer is 5cm when thermometer is in contact with water at its triple point , calculate the measured temperature when the mercury column is 6cm and find the length of the mercury column when it is at steam point



1
Expert's answer
2021-02-04T14:40:53-0500

Answer

According to given data

1cm=0.01/5cm=0.002°C

So temperature

T=0.002"\times" 6=0.012°C

Now length

L=100/5=20cm